OpenStudy (anonymous):

What are the types of roots of the equation below? x 4 - 81 = 0 Four complex Two complex and two real Four real

OpenStudy (anonymous):

four real

OpenStudy (amoodarya):

\[x^4-81=0\\(x^2-9)(x^2+9)=0\\x^2=9 ,x^2=-9\\ \rightarrow x^2=9 \\x=\pm3\\x^=-9 \rightarrow x=\pm 3i\]

OpenStudy (gorv):

(x^2-9)(x^2+9)=0

OpenStudy (amoodarya):

2 real and 2 complex